the sum of 2positive integer is 98. theinteger in ration of 3:4 find integers
Answer by malaydassharma(59)

Let the integers x ad y. Hence,
x+y = 98 ....(i)
x:y=3:4
x=3y/4 .....(ii)

From (i) and (ii) , we get

(3y/4)+y=98
Or, 7y= 4X98
Or, y = (4X98)/7 = 56
y = 56
x = (3X56)/4 = 42